# Masked Training for Robust Arrhythmia Detection from Digitalized Multiple Layout ECG Images

### v1

@@ ABSTRACT
Electrocardiogram (ECG) as an important tool for diagnosing cardiovascular diseases such as arrhythmia. Due to the differences in ECG layouts used by different hospitals, the digitized signals exhibit asynchronous lead time and partial blackout loss, which poses a serious challenge to existing models. To address this challenge, the study introduced PatchECG, a framework for adaptive variable block count missing representation learning based on a masking training strategy, which automatically focuses on key patches with collaborative dependencies between leads, thereby achieving key recognition of arrhythmia in ECGs with different layouts. Experiments were conducted on the PTB-XL dataset and 21388 asynchronous ECG images generated using ECG image kit tool, using the 23 Subclasses as labels. The proposed method demonstrated strong robustness under different layouts, with average Area Under the Receiver Operating Characteristic Curve (AUROC) of 0.835 and remained stable (unchanged with layout changes). In external validation based on 400 real ECG images data from Chaoyang Hospital, the AUROC for atrial fibrillation diagnosis reached 0.778; On 12 x 1 layout ECGs, AUROC reaches 0.893. This result is superior to various classic interpolation and baseline methods, and compared to the current optimal large-scale pre-training model ECGFounder, it has improved by 0.111 and 0.19.
KEYWORDS
Deep Learning, Masked Training, Multiple Layout ECG Images, Digitalized ECG
@@ INTRODUCTION
Cardiovascular disease (CVD) is one of the world’s deadliest disease, causing approximately 17.8 million deaths according to 2017 statistics 1. Electrocardiogram (ECG) is a key tool for the diagnosis and treatment of CVD, revealing important pathophysiological information by recording the electrical activity of the heart. In recent years, deep learning techniques have been shown to rival trained doctors in ECG signal and ECG arrhythmia detection tasks. Despite recent advances in digital ECG equipment, the use of photographs, printed screens, or paper ECGs is still prevalent in the Global South, where ECG slices are far less expensive to store than signal data, and where these recordings not only contain a rich history of CVD, but also reflect its diversity. In underdeveloped regions, this non-digitized data format highlights the urgent need for high-quality digitization and automated analysis tools 2. Thus, digitizing traditional ECG images and building algorithms capable of automated diagnosis based on them will both help to deepen our understanding of CVD and, hopefully, significantly improve the diagnosis and treatment of underrepresented and medically under-resourced populations.
@@
ECG images exist in a variety of layouts. For example, in Fig. 1(a), different forms such as common 3 x 4 + II layout, 6 x 2 layout or complete 12-lead may be presented. In existing studies, ECG images with different layouts are usually directly inputted into an image classification model for classification3, 4. Data augmentation (e.g., rotation, folding, light and shadow adjustments, disrupting lead positions, etc.) of such images can effectively improve the robustness of the model so that it still maintains a high accuracy in predicting real-world data. Data augmentation (e.g., rotation, folding, light and shadow adjustments, disrupting lead positions, etc.) of such images can effectively improve the robustness of the model so that it still maintains a high accuracy in predicting real-world data. However, image classification models are prone to focus excessively on the image background rather than the ECG signal itself. Although ECG all-in-one 4 provides more accurate interpretability by calculating the classifier gradient through Jacobi matrix sensitivity, its interpretation results still focus on the image background. This tends to raise questions from experts about its interpretability, as physicians have more confidence in the prediction results based on the ECG signal itself than the image model.
With the development of methods for digitizing ECG images, ECG image data can be accurately converted into ECG signal data5, 6, making it possible to use the signal data for subsequent analysis. As shown in Fig. 1(b), the digitization of ECG image with different layouts of image produces 2.5s asynchronous, 5s asynchronous and synchronous signals 7 and “Partial Blackout” which covers situations where a variable number of features become unavailable for some time8. This digitized signal ECG presents two major challenges:
@@
  • We test on ECG images collected at Chaoyang Hospital. The images were converted into signal data by digitization methods and based on which external validation was performed. The results show that our method is effective in capturing critical patches and making accurate diagnosis. In particular, on the external validation cohort of AF ECG images from Chaoyang Hospital, the model evaluation achieved an AUROC of 0.778, which effectively demonstrates the ability of the method to identify the key patches and achieve an accurate diagnosis of AF.
RESULTS
To comprehensively evaluate the performance of our proposed PatchECG, we conducted a series of rigorous experiments and we conducted a series of rigorous experiments and compared it with several key baseline models, including: 1) The typical baseline 9 in the main follow MissTSM paper uses zero padding, KNN interpolation, and SAITS interpolation10, respectively, and combines them with SimMTM for prediction 11; 2) Medformer is a model specifically designed for biosignals, which introduces six different noise perturbations intra- and inter-granularity during the training phase, thus enhancing robustness to missing 12;3) TimeXer, on the other hand, provides auxiliary cues to the model by explicitly modeling missing patterns and treating the missing as a kind of external information13. We explored the diagnostic accuracy of the generated signal data digitized from ECG images by different layouts of signal data and the signal data digitized from different layouts of ECG images from the real Chaoyang Hospital.
PatchECG Achieve Better Results on Different Layouts
Figure 2: Results of PatchECG and Other Methods with Different Layouts
PatchECG is Highly Scalable
To effectively utilize existing excellent models, we explored three different modules for encoding patches. We primarily adopted representative models such as Net1D, ECGFounder, and Projection as PatchEncoders: 1. Net1D: a residual one-dimensional convolutional model combined with an SE module 14. 2. ECGFounder: the latest foundational model pre-trained on 10 million ECGs; we froze its parameters and used it as a PatchEncoder to encode patches 15. 3. Projection: inspired by PatchTST, it performs simple projection on patches through a fully connected layer 16.
Table 1: Different modules as PatchEncoder
Method
3 x 4
3 x 4 + II
3 x 4 + II + V1
6 x 2
6 x 2 + II
12 x 1
Random
Ours-Resnet 0.831 $\pm$ 0.019 0.829 $\pm$ 0.019 0.830 $\pm$ 0.022 0.833 $\pm$ 0.022 0.832 $\pm$ 0.022 0.829 $\pm$ 0.026 0.835 $\pm$ 0.022
We adopted the same training, validation, and test set partitions as in Experiment 1 and conducted 5-fold cross-validation. As shown in Table 1, due to the advantages of the method, we can use different modules as PatchEncoders. When dealing with multi-label classification tasks under different layouts, the model effectively captures the asynchronous signal dependencies caused by layouts differences. Therefore, even if the input signal layouts vary, the model can still achieve a close average AUROC across different layouts. A detailed analysis of the performance of each PatchEncoder module is as follows: Projection: Due to its simple structure, it is difficult to learn complex information about patches, making it the worst performing as a PatchEncoder. Net1D: Trained from scratch, it achieved a more robust average AUROC performance in testing after 5-fold cross-validation, achieving the second-best result. Large-scale ECG pre-trained model (ECGFounder): Its parameters were frozen and directly used for patch encoding, achieving the best average AUROC. In summary, our method can effectively utilize existing excellent models (including large-scale pre-trained models) as PatchEncoder, combined with an adaptive variable block number missing mechanism, enabling PatchECG to maintain high accuracy when processing both asynchronous and synchronous ECG data.
PatchECG Achieve Better Results on Generating ECG Images
To further verify the adaptability of the model in complex real-world scenarios, the ECG-image-kit tool was used to generate ECG images. During the image generation process, common practical interference factors were simulated, and noises such as rotation, wrinkling, and light and shadow changes were added to more realistically reproduce images in clinical settings. Due to limitations of the tool itself, the generated ECGs adopted a 3 × 4 layout. Subsequently, the generated ECG images were subjected to signal reconstruction using the digital method that won first place in the George B. Moody PhysioNet Challenges 2024 5. The digitized signals were used as inputs for model testing, and the labels used in the testing phase were consistent with those used in the training phase. We adopted the AUROC as the evaluation metric and used 5-fold cross-validation to improve the stability and reliability of the results. The experimental results are shown in Figure 3.
Figure 3: Digitally generated ECG image
PatchECG Achieve Better Results on Real Hospital ECG Images
On the external test set of Chaoyang Hospital, which mainly consists of 12 x 1 layout, with some being 3 x 4 + II and 6 x 2 layout, our method achieved the best AUROC among all AF baseline predictions (as shown in Table 2). The overall performance of ECGFounder’s predictions is good. Although some methods have higher recall rates, their specificity is significantly lower, making them prone to misclassifying negative samples as positive, leading to an increased risk of misdiagnosis. Due to pre-training on tens of millions of ECGs, ECGFounder exhibits strong robustness. When there is a certain degree of data loss, the method fills in the data to a complete 12 x 1 layout through zero-padding before inputting it into the model. On the external AF cohort of Chaoyang Hospital, the three imputation strategies performed similarly in terms of AUROC scores, with almost no significant differences. This indicates that imputation does not significantly improve the model’s accuracy. Medformer performed the best among all baseline methods. This method introduces six data augmentation strategies during the training stage, which act on its internal and external granularity modeling modules, respectively, to enhance the model’s robustness to noise and missing data. When dealing with the Chaoyang Hospital cohort with different layouts and poor digitization quality, Medformer achieved an AUROC of 0.605, demonstrating strong adaptability and stability. TimeXer performed well on high-quality digitized signal data, but its performance significantly decreased in the Chaoyang Hospital cohort with poor digitization quality. This method treats missing data as exogenous information and uses it to assist in the modeling and prediction of endogenous information. However, when digitization is poor, not only is the endogenous information itself inaccurate, but the generated missing pattern information also becomes inaccurate. The exogenous information originally used to guide the model becomes a source of interference instead. The introduction of this noise weakens the model’s judgment ability, leading to unsatisfactory prediction results on real Chaoyang Hospital image data.
Method Accuracy Precision Recall Specificity F1 AUROC
Ours 0.745 0.743 0.750 0.740 0.746 0.778
ECGFounder $0.625$ $0.605$ $0.720$ $0.530$ $0.658$ $0.667$
SimMTM $0.463$ $0.332$ $0.600$ $0.395$ $0.427$ $0.541$
SimMTM-KNN $0.485$ $0.480$ $0.540$ $0.430$ $0.512$ $0.539$
SimMTM-SAITS $0.490$ $0.491$ $0.550$ $0.430$ $0.519$ $0.541$
Medformer $0.498$ $0.498$ $0.640$ $0.335$ $0.560$ $0.605$
TimeXer $0.489$ $0.487$ $0.460$ $0.515$ $0.473$ $0.485$
Method Accuracy Precision Recall Specificity F1 AUROC
Ours 0.833 0.723 $0.810$ 0.845 0.764 0.893
ECGFounder $0.573$ $0.430$ 0.860 $0.430$ $0.573$ $0.703$
SimMTM $0.485$ $0.486$ $0.525$ $0.445$ $0.505$ $0.450$
SimMTM-KNN
SimMTM-SAITS
Medformer $0.450$ $0.332$ $0.640$ $0.355$ $0.437$ $0.510$
TimeXer $0.473$ $0.289$ $0.390$ $0.510$ $0.331$ $0.423$
Based on this, we further selected samples arranged in a 12 × 1 layout for additional evaluation (as shown in Table 3). Our method still achieved the best performance in terms of the AUROC metric, demonstrating strong robustness. ECGFounder achieved the highest recall rate in this 12 × 1 layout evaluation, exhibiting a strong ability in identifying AF patients and more accurately detecting positive samples. However, its specificity is relatively low, making it prone to misdiagnosing normal individuals as AF patients, posing a certain risk of overdiagnosis. The performance of MedFormer improves with the integrity of lead data, showing significant improvement in the complete 12 × 1 layout, with a recall rate reaching 0.64. In contrast, the other baseline methods largely fail in scenarios with poor digital quality, struggling to effectively model or make accurate predictions. This further highlights the adaptability and advantages of our method in real complex environments.
Case Study
DISCUSSION
Based on the aforementioned concerns, we propose PatchECG, a novel predictive architecture that addresses arbitrarily layout signal inputs. It combines a masking training strategy and a patch-oriented attention mechanism to achieve intelligent recognition of arbitrarily layouts ECG arrhythmias. We systematically evaluated the proposed method on three types of data, including ECG data with different lead layouts, ECG image signals generated through digitization, and real clinical ECG images from Chaoyang Hospital. In all test scenarios, the model exhibited good predictive performance. Specifically, when training with raw signal data and testing on signal data after image digitization, the model still achieved stable and accurate prediction results7. In the external test set from Chaoyang Hospital, which had relatively poor quality, our method achieved an AUC of 0.778 in the AF prediction task, demonstrating strong robustness. Furthermore, when evaluated on additionally filtered synchronous 10-second ECG data, the model’s AUC increased to 0.893. The interpretability of the model is of great significance for enhancing clinicians’ understanding and trust in AI prediction results.
Data usage and test setup, we use PTB-XL data as the training data source in this study. There are certain limitations in data usage, which to some extent restrict the model’s generalization ability to broader patient populations and more complex pathological types. During the testing phase, we utilize the ECG-image-kit tool to generate ECG images and digitize them to evaluate the model. However, due to limitations in digitization methods, only 3 x 4 layout can be accurately processed, and other layout types of images cannot be effectively digitized. Therefore, in the generated images, we primarily focus on the 3 x 4 layout. In recent years, various digitization methods have emerged, but solutions that truly efficiently and accurately restore high-quality signals remain limited. It is particularly noteworthy that current digitization methods generally lack effective mechanisms to deal with lead adhesion. When lead adhesion occurs, the signals in the image overlap, which can easily be misjudged by the model as pathological features such as ST segment deviation, leading to deviations in prediction results. Therefore, how to improve the digitization quality of ECG images, especially developing digitization algorithms capable of identifying and correcting lead adhesion issues, is one of the important directions for achieving higher-quality AI-ECG diagnosis research in the future.
CONCLUSION
@@ METHODS
Artificial intelligence has made significant progress in the application of ECG signal arrhythmia detection. Most jobs typically use synchronized 10-s complete ECG signal data, such as using CNNsau2025comparison, chen2024congenital,sangha2022automated, CNN + RNN17, CNN + Transformer18 and a hybrid model combining CNN and self-attention mechanism 15, etc. When the input data dimensions are inconsistent, interpolation is usually used to uniformly fill the data to meet the input requirements of the model. Common interpolation strategies include zero padding, traditional machine learning interpolation methods, and deep learning based interpolation methods. SimMTM11, Medformer12, S4D-ECG19, CNN + SE15 and methods such as BiLSTM 20 use zero padding strategy to complete missing data before making predictions. These methods can effectively capture the connections between leads and model time-dependent features in synchronizing ECG signal data. Among them, Medformer 12 introduces six data augmentation strategies to make it robust in the face of a certain degree of missing data. In addition, a Random Lead Masking (RLM) training strategy was proposed to ensure stable performance of the model even in the missing of complete leads. By randomly masking a signal segment of about 1.5 seconds, the local loss caused by electrode detachment is simulated to improve the adaptability of the model to small-scale block like loss 21. Using methods such as KNN and SAITS to interpolate missing data can to some extent complete data completion, allowing for subsequent prediction using existing synchronous 10s ECG signal analysis methods 9.
However, the above methods all rely on interpolation and modeling strategies. Even with the use of the most advanced interpolation techniques, interpolation errors will still propagate in subsequent modeling, limiting the improvement of model performance. Meanwhile, the uncertainty introduced by interpolation can also reduce the interpretability of prediction results, posing potential challenges for clinical applications.
@@
MissTSM 9 proposed a new embedding scheme called Time Feature Independent (TFI) embedding, which considers each combination of time steps and features (or channels) as independent labels and encodes them into high-dimensional space, effectively capturing spatiotemporal dependencies without interpolation and achieving direct modeling of missing data. TimeXer introduces external information modeling to enable standard Transformers to coordinate the processing of relationships between internal and external sources 13. GRU-D proposed a GRU model with attenuation mechanism, which considers missing signals as external signals to enhance the modeling ability of the model on the original time series 22. BiTGraph further designed a biased temporal convolutional graph neural network to jointly model temporal dependencies and spatial structural features23. The above methods can directly model based on missing patterns and effectively capture potential temporal and multivariate dependencies.
@@
In summary, although progress has been made in ECG signal data imputation and direct modeling based on missing data, so far there has been no research attempting to solve the asynchronous signal and "Partial Blackout" missing problems caused by different layouts of ECGs. Therefore, this study proposes PatchECG, which designs a novel and universal architecture that adapts to signal data of different layouts, arbitrary lead numbers and lengths without introducing additional noise, and provides interpretability analysis of key patches. This method provides new insights for predicting CVD based on ECGs.
@@
In order to enable the model to effectively learn different layouts of ECG, we propose a masking training strategy to improve the robustness of the model. Before the training starts, mask the ECG data by assigning random starting points, initial positions, and lengths to each lead, and obtain the masked ECG data ${\textstyle\sum_{i=1}^{C}}n_{i}$.
As shown in part a of the Figure 5, first we perform block processing on the data after the mask, and perform equidistant segmentation of the ECG data without coverage, using a window size of $P$. After segmentation, we obtain $N$ patches, $N={\textstyle\sum_{i=1}^{C}}n_{i}$. Among them:
$n_{i}=\left\lfloor\frac{c_{i}}{P}\right\rfloor,C\in\left\{c_{i}\mid i=1,2\cdots C\right\}$ (1)
$c_{i}$ represents the $i$ th lead, $n_{i}$represents the number of $i$ th lead blocks, and the last incomplete part is discarded. After cutting, it is divided into three main types of patches: complete patches, partially missing patches, and completely missing patches.
$m=\left\{M_{i,j}\in\mathbb{R}^{P},i=1,2,\dots,C,j=1,2,\dots,\left\lfloor\frac{c}{P}\right\rfloor\right\}$ (2)
$m$ represents each block. Add internal markers to each block, generate a marker with the same dimension as the block size based on the missing blocks, with a corresponding value of 1 for existing signals and 0 for missing signals, and perform concatenation, gets $\tilde{m}$
$\tilde{m}=\left\{{\tilde{M}_{i,j}}\in\mathbb{R}^{2\times P},i=1,2,\dots,C,j=1,2,\dots,\left\lfloor\frac{c}{P}\right\rfloor\right\}$ (3)
Input into PatchEncoder to obtain $\bar{m}\in{\mathbb{R}^{D}}$, where $D$ represents the dimension of the embedding layer encoded and output by PatchEncoder for each block. The existing signal based methods include numerous excellent models, which typically take fixed layout and fixed length signals as inputs. We effectively utilize these models and transform them into PatchEncoders to encode the input in patch form. In this way, PatchECG can process ECG signal data with any number of leads, any length, and any layout. We mainly use SE block with attention mechanism as the basic module, combined with Net1d and residual structure to extract high-dimensional features14.

### v3

@@ ABSTRACT
Background
Electrocardiograms are indispensable for diagnosing cardiovascular diseases, yet in many clinical and resource-limited settings they exist only as paper printouts or photographs stored in multiple recording layouts. Converting these images into digital signals introduces two key challenges: temporal asynchrony among leads and partial blackout missing, where contiguous signal segments become entirely unavailable. Existing models cannot adequately handle these concurrent problems while maintaining interpretability.
Methods
[…]
Results
[…]
Conclusions
PatchECG provides a robust, interpolation-free, and interpretable solution for arrhythmia detection from digitized ECG images across diverse layouts. Its direct modeling of asynchronous and partially missing signals, combined with clinically aligned patch-level attention, positions it as a practical tool for cardiac diagnostics from legacy ECG image archives in real-world clinical environments.
Plain language summary
ECG recordings of the heart’s electrical activity are a cornerstone of cardiac diagnosis. In many hospitals worldwide, ECGs exist only as paper printouts or photographs taken in different layouts. Converting these images into digital signals introduces missing and misaligned data that existing models cannot reliably handle. We developed PatchECG, a model that works directly with incomplete and misaligned ECG signals across any recording format, without artificially filling in the gaps. Instead of patching over missing data, PatchECG learns to focus on the signal segments that matter most—the same features experienced cardiologists examine. Validated on real hospital images, PatchECG consistently outperformed existing methods and produced interpretable, clinically meaningful outputs. This work offers a practical pathway to unlock decades of paper ECG archives for automated cardiac diagnosis, with particular relevance for under-resourced healthcare settings.
@@ Introduction
Cardiovascular disease (CVD) is one of the world’s deadliest disease, causing approximately 17.8 million deaths according to 2017 statistics 21. Electrocardiogram (ECG) is a key tool for the diagnosis and treatment of CVD, revealing important pathophysiological information by recording the electrical activity of the heart. In recent years, deep learning techniques have been shown to rival trained doctors in ECG signal and ECG arrhythmia detection tasks. Despite recent advances in digital ECG equipment, the use of photographs, printed screens, or paper ECGs is still prevalent in the Global South, where ECG slices are far less expensive to store than signal data, and where these recordings not only contain a rich history of CVD, but also reflect its diversity. In underdeveloped regions, this non-digitized data format highlights the urgent need for high-quality digitization and automated analysis tools 10. Thus, digitizing traditional ECG images and building algorithms capable of automated diagnosis based on them will both help to deepen our understanding of CVD and, hopefully, significantly improve the diagnosis and treatment of underrepresented and medically under-resourced populations.
In clinical practice, when AI models are applied to ECG image data, image-based classification methods suffer from a fundamental interpretability limitation: their predictions are often driven by background features rather than ECG waveform morphology, which undermines clinician trust. In contrast, signal-based models can provide waveform-level explanations that directly correspond to established diagnostic criteria (e.g., P-wave morphology, RR interval regularity). By digitizing ECG images into signals and applying PatchECG, clinicians gain access to interpretable, signal-level diagnostic support from legacy image archives—particularly valuable in settings where only paper or photographic ECG records exist.
@@
ECG images exist in a variety of layouts. For example, in Fig. 1(a), different forms such as common 3 x 4 + II layout, 6 x 2 layout or complete 12-lead may be presented. In existing studies, ECG images with different layouts are usually directly inputted into an image classification model for classification24, 9, 33. Data augmentation (e.g., rotation, folding, light and shadow adjustments, disrupting lead positions, etc.) of such images can effectively improve the robustness of the model so that it still maintains a high accuracy in predicting real-world data. However, image classification models are prone to focus excessively on the image background rather than the ECG signal itself. Although ECG all-in-one 9 provides more accurate interpretability by calculating the classifier gradient through Jacobi matrix sensitivity, its interpretation results still focus on the image background. This tends to raise questions from experts about its interpretability, as physicians have more confidence in the prediction results based on the ECG signal itself than the image model.
With the development of methods for digitizing ECG images, ECG image data can be accurately converted into ECG signal data17, 8, making it possible to use the signal data for subsequent analysis. As shown in Fig. 1(b), the digitization of ECG image with different layouts of image produces 2.5s asynchronous, 5s asynchronous and synchronous signals 25 and “Partial Blackout” which covers situations where a variable number of features become unavailable for some time14. This digitized signal ECG presents two major challenges:
@@
[…]
@@ Methods
Artificial intelligence has made significant progress in the application of ECG signal arrhythmia detection. Most works typically use synchronized 10-s complete ECG signal data, such as using CNN25, 3, 24, CNN+RNN29, CNN+Transformer37 and hybrid models combining CNN and self-attention mechanism18, etc. In recent years, Transformer-based methods have demonstrated outstanding performance in ECG classification. MSGformer16 fuses multi-lead spatial features through self-attention and introduces multi-scale grid attention to capture temporal features at different scales. ECGTransForm7 combines multi-scale convolutions with a bidirectional Transformer to simultaneously capture spatial features and bidirectional temporal dependencies. AC-ITCT15 introduces a Temporal Convolutional Transformer with adaptive temporal resolution adjustment, balancing classification accuracy and interpretability. MDOT36 converts one-dimensional ECG signals into two-dimensional oscillographic representations via an OSC module, incorporating clinical electrophysiological indicators and achieving lightweight deployment through momentum distillation. Chatterjee et al.1 utilized 8.2 million unlabelled ECGs for masked patch modeling pre-training, building a 1D Vision Transformer based on self-supervised learning. Yang et al.34 systematically analyzed the impact of lead combinations on classification performance from a medical anatomical perspective, proposing a five-lead-grouping strategy and L5G-Net architecture. When the input data dimensions are inconsistent, interpolation is usually used to uniformly fill the data to meet the input requirements of the model. Common interpolation strategies include zero padding, traditional machine learning interpolation methods, and deep learning based interpolation methods. SimMTM5, Medformer31, S4D-ECG13, CNN+SE18 and methods such as BiLSTM20 use zero padding strategy to complete missing data before making predictions. These methods can effectively capture the connections between leads and model time-dependent features in synchronizing ECG signal data. Among them, Medformer31 introduces six data augmentation strategies to make it robust in the face of a certain degree of missing data. In addition, a Random Lead Masking (RLM) training strategy was proposed to ensure stable performance of the model even in the missing of complete leads20, 3, 38. By randomly masking a signal segment of about 1.5 seconds, the local loss caused by electrode detachment is simulated to improve the adaptability of the model to small-scale block like loss35. Using methods such as KNN and SAITS to interpolate missing data can to some extent complete data completion, allowing for subsequent prediction using existing synchronous 10s ECG signal analysis methods22.
However, the above methods all rely on interpolation and modeling strategies. Even with the use of the most advanced interpolation techniques, interpolation errors will still propagate in subsequent modeling, limiting the improvement of model performance. Meanwhile, the uncertainty introduced by interpolation can also reduce the interpretability of prediction results, posing potential challenges for clinical applications. Moreover, the aforementioned Transformer methods all assume synchronized 10-second complete signals. When facing asynchronous 2.5s or 5s ECG data caused by different layouts, they cannot adapt to the temporal inconsistency among leads and large-scale Partial Blackout missing. For example, MSGformer relies on fixed-length time window partitioning, ECGTransForm requires complete temporal sequence context, MDOT’s oscillographic conversion requires temporal alignment, and the PatchECG by Chatterjee et al.1, although adopting a patch strategy, focuses on large-scale self-supervised pre-training without specifically modeling the Partial Blackout missing in asynchronous ECG.
@@
MissTSM 22 proposed a new embedding scheme called Time Feature Independent (TFI) embedding, which considers each combination of time steps and features (or channels) as independent labels and encodes them into high-dimensional space, effectively capturing spatiotemporal dependencies without interpolation and achieving direct modeling of missing data. TimeXer introduces external information modeling to enable standard Transformers to coordinate the processing of relationships between internal and external sources 32. GRU-D proposed a GRU model with attenuation mechanism, which considers missing signals as external signals to enhance the modeling ability of the model on the original time series 2. BiTGraph further designed a biased temporal convolutional graph neural network to jointly model temporal dependencies and spatial structural features4. The above methods can directly model based on missing patterns and effectively capture potential temporal and multivariate dependencies.
@@
Multi-Layout ECG Processing Methods
In recent years, research on multi-layout ECG processing has begun to emerge. Wang and Yang30 proposed a paper ECG digitization and classification framework combining adaptive rotated convolution with a state space augmented Transformer, capable of handling paper ECG images with rotations, creases, and high-level noise, but its classification model still predicts based on synchronized signals after digitization without modeling asynchronous missing. Yang et al.34 analyzed the impact of lead combinations on classification performance from a medical anatomical perspective, proposing a five-lead-grouping strategy that reveals the importance of spatial complementarity among leads. These works provide important references for multi-layout ECG processing and understanding lead spatial relationships, but none address the signal asynchrony and large-scale Partial Blackout missing caused by different ECG layouts.
In summary, although progress has been made in ECG signal data imputation, direct modeling based on missing data, and multi-layout ECG processing, so far there has been no research attempting to solve the asynchronous signal and “Partial Blackout” missing problems caused by different layouts of ECGs. Existing Transformer methods16, 7, 15, 36 all assume temporally aligned and complete signals; the PatchECG by Chatterjee et al.1 focuses on large-scale self-supervised pre-training without designing specific mechanisms for asynchronous ECG; and the paper ECG digitization method30 only addresses the conversion from images to signals. Therefore, this study proposes PatchECG, which designs a novel and universal architecture that adapts to signal data of different layouts, arbitrary lead numbers and lengths without introducing additional noise, and provides interpretability analysis of key patches. Unlike existing methods, PatchECG directly models the Partial Blackout missing in asynchronous ECG through an adaptive variable block number missing learning mechanism and a patch-based collaborative dependency attention mechanism, effectively capturing spatiotemporal dependencies across leads without interpolation, providing new insights for predicting cardiovascular diseases based on ECGs with different layouts.
@@
In order to enable the model to effectively learn different layouts of ECG, we propose a masking training strategy to improve the robustness of the model. Before training starts, each lead is independently masked by randomly sampling a starting position and a mask length. Specifically, for the $i$-th lead of total length $L$ (in sampling points), the mask starting position $st_{i}$ and mask length $l_{i}$ are independently drawn from uniform distributions:
$st_{i}\sim\mathcal{U}(0,\,L),\quad l_{i}\sim\mathcal{U}(st_{i},\,L-st_{i})$ (1)
ensuring the masked region $st_{i},\,st_{i}+l_{i})$ lies entirely within the signal boundary. Each lead contains at most one contiguous masked segment. All masked signal values are set to NaN to distinguish them from observed zero-valued signal points.
As shown in part a of the Figure 2, after masking, we perform equidistant non-overlapping segmentation on each lead using a fixed window size of $P$. For the $i$-th lead, the number of complete patches $n_{i}$ retained is:
$n_{i}=\left\lfloor\frac{c_{i}}{P}\right\rfloor,\quad C\in\left\{c_{i}\mid i=1,2\cdots C\right\}$ (2)
where $c_{i}$ denotes the total length of the $i$-th lead. The floor operation $\lfloor\cdot\rfloor$ ensures that only complete patches of exactly $P$ points are retained: when the remainder $r_{i}=c_{i}\bmod P>0$, the last $r_{i}$ points are discarded without zero-padding or interpolation, so that no artificial values are introduced. The total patch count across all leads is $N=\sum_{i=1}^{C}n_{i}$.After segmentation, each patch $m$ is defined as:
$m=\left\{M_{i,j}\in\mathbb{R}^{P},\ i=1,\dots,C,\ j=1,\dots,\left\lfloor\frac{c_{i}}{P}\right\rfloor\right\}$ (3)
Here $m$ denotes an individual patch. For each patch, we construct a binary mask indicator of the same length $P$, assigning a value of 1 to observed signal points and 0 to masked (missing) points. This indicator is then concatenated with the patch along the feature dimension to form the augmented patch $\tilde{m}$ is :
$\tilde{m}=\left\{{\tilde{M}_{i,j}}\in\mathbb{R}^{2\times P},\ i=1,\dots,C,\ j=1,\dots,\left\lfloor\frac{c_{i}}{P}\right\rfloor\right\}$ (4)
Based on the sum of indicator values $\sum_{k}\mathbf{b}_{i,j}[k]$, each patch is classified into one of three types: a complete patch (all $P$ values observed), a partially missing patch ($0<\sum_{k}\mathbf{b}_{i,j}[k]<P$), or a completely missing patch ($\sum_{k}\mathbf{b}_{i,j}[k]=0$).
Input into PatchEncoder to obtain $\bar{m}\in\mathbb{R}^{D}$, where $D$ represents the embedding dimension output by PatchEncoder for each block. The existing signal-based methods include numerous excellent models that typically take fixed-layout and fixed-length signals as inputs. We effectively utilize these models by transforming them into PatchEncoders to encode inputs in patch form. In this way, PatchECG can process ECG signal data with any number of leads, any length, and any layout. We primarily employ the SE block with an attention mechanism as the basic module, combined with Net1D and a residual structure to extract high-dimensional features 12.